Adult tickets for a play cost $11 and child tickets cost $10. If there were 23 people at a perfromance and the theater collected $246 from tickets sales, how mant childern attended the play?
Answer by checkley71(8403)

A+C=23 OR A=23-C FOR THE NUMBER OF TICKETS SOLD.
11A+10C=246
NOW REPLACE A WITH(23-C) IN THE SECOND EQUATION & SOLVE FOR C
11(23-C)+10C=246
253-11C+10C=246
-C=246-253
-C=-7
C=7 CHILDREN'S TICKETS WERE SOLD. (NO IDEA HOW MANY CHILDREN ATTENDED)
23-7=16 ADULT TICKETS SOLD.
PROOF
11*16+10*7=246
176+70=246
246=246